OpenStudy (anonymous):

Ok I just need to know how to set this problem up and then I can solve it. One canned juice drink is 15% orange juice; another is 5% orange juice. How many litres of each should be mixed together in order to get 10L that is 7 % orange juice. I have to use a system of equations to solve this. Does x+y =7 and then 15x + 5y = (10)(7) make sense?

OpenStudy (anonymous):

I would just use one variable for OJ, x, and make it more like: A(0.15x) + B(0.05x) = 10(0.07x) ; but this might require a little diff eq, if not then linear algebra. Does this work for you?

OpenStudy (anonymous):

15x+5(10-x)=7*10

OpenStudy (anonymous):

why 5(10 - x)?

OpenStudy (anonymous):

DoinHomework that makes sense but I still need another equation so that I can solve the system. But even if I do it the way you showed I still have to multiply the equation by 10 so I do not have to mess with the decimals.

OpenStudy (anonymous):

sorry 100 not ten